OpenStudy (anonymous):

Solve: 49x2 - 140x = -100

OpenStudy (anonymous):

the equation is the same as 49x^2 - 140x +100 = 0 represent 140 as the sum of 2 numbers whose product is equal to the product of the co-efficient of x^2 and the constant

OpenStudy (anonymous):

i am still confused so the answer is 0

OpenStudy (anonymous):

Okay .. consider 140x = 70x +70x then the equation becomes 49x^2 - 70x - 70x + 100 = 0 factorize the first two and the second two terms to get the equation in the form of (x + a)(x + b)=0 then the values of X can be got

OpenStudy (anonymous):

got it .......i am pretty sure i might be back

OpenStudy (anonymous):

so the answer is -7/10, 7/10

OpenStudy (anonymous):

its +7/10 , +7/10 :)

OpenStudy (anonymous):

hero

OpenStudy (anonymous):

:)